# SelectLLM: Query-Aware Efficient Selection Algorithm for Large Language Models

### v1

@@ 5.1 Baseline Models
The following baseline models are included for comparison:
@@
    Random: This represents the mean performance of uniformly randomly selecting an LLM subset from all possible ($12^{k}-$ 1) subsets for each query. We report mean evaluation scores across 1,000 independent runs to avoid biases.
@@
    Individual Models: This is the performance of individual models with maj@10.
@@
    All LLMs: This baseline reports the mean accuracy of maj@(10×$\times|L|$) based on the combined pool of 10 generations from each LLM and is similar to the approach of Li et al. (2024).
@@
  6. 6.
    Top-s LLMs: For this baseline, we consider the responses of the $s$ top-scoring LLMs for the majority-vote. We determine the top performing models by the aggregate accuracies of each model on the test sets of corresponding datasets.
For All LLMs, LLM-Blender and Top-s LLMs, the latency remains constant since it necessitates inference with all LLMs to determine the performance.
@@ 6 Results and Discussion
LLM-Blender Jiang et al. (2023) 75.28 19.00 60.27 16.40
Top-s LLMs 77.48 (s=4) 19.00 65.75 (s=2) 16.40
gemma-7b-lm 71.27 7.10 63.73 3.00
mistral-7b-lm 60.50 3.70 61.57 1.80
metamath-7b-lm 67.25 4.70 41.76 2.40
llama2-7b-lm 48.10 2.30
llama2-13b-chat 49.20 1.80 52.94 4.80
mistral-7b-it 56.71 1.00 53.92 1.10
gemma-7b-it 42.23 0.70 50.72 1.00
SelectLLM MLC + LabelledMaxConf 75.66 (s=4) 14.69 65.68 (s=2) 4.78
MLC + MaxConf 77.48 (s=4) 16.50 65.68 (s=2) 4.78
MLC + WeightedMaxConf 77.94 (s=4) 16.50 65.81 (s=2) 4.78
[…]
@@
Table 2 presents the performance results for the oracle, baselines, and the proposed SelectLLM models across both GSM8K and MMLU datasets. We have also reported respective inference latencies to analyze the efficiency of different models. We make the following major observations.
Performance of Individual Models: Among the considered LLMs, metamath-7b-lm, a specialized mathematical LLM, emerges as the best performing model for the mathematical reasoning task (GSM8K) but performs worse on the natural language understanding/reasoning task (MMLU). This indicates the LLM diversity and the need for the effective combination of their capabilities. Gemma-7b-lm performs the best on MMLU and second best on the GSM8K data. Chat LLMs perform poorly on GSM8K; however, for MMLU, their performance is mixed; some non-chat LLMs achieve better scores. The latency for non-chat models is expected to be higher than for chat models, as they utilize a 5-shot COT, whereas chat models employ a 0-shot COT strategy.
Performance of Baselines: We have tested four baseline models (Random, All LLMs, LLM Bender, and Top-s LLMs) to understand the effect of ensembling LLMs. It can be observed that the performance of the Random baseline model is better than that of many individual LLMs, demonstrating the promise of utilizing multiple LLMs. To our surprise, including all LLMs does not perform better than selecting only the top few. This contradicts the findings of Li et al. (2024) and can be attributed to two reasons: (i) the authors have used only a few LLMs, and (ii) they have employed very powerful LLMs, leaving out weaker LLMs. The Top-s LLMs model emerges as the best baseline, showing the promise of selecting the best-performing LLMs. However, the latency of all these baselines is high (except Random baseline where accuracy is lower), making them impractical for real-world usage.
@@
The Effect of Different Policies with the SelectLLM Algorithm: It can be observed that with the policy LabelledMaxConf, the performance on the GSM8K dataset is the lowest. This may be because the policy relies on both the MLC prediction and the confidence scores, where the classifier almost always predicts metamath-7b-lm as the class label (because metamath-7b-lm is a specialized model for math, and 88% of the SLData training has this label). However, this impact is negligible for MMLU, where the label distribution for all LLMs is not as highly skewed. This limitation is overcome by MaxConf and WeightedMaxConf, which relax the constraint on MLC label prediction and only operate on confidence scores. This allows models to incorporate other LLMs and push the performance, particularly for the GSM8K data. Mathematical tuning in WeightedMaxConf allows policies to select LLMs more effectively and improve the scores. Overall, the WeightedMaxConf policy emerged as the best performing, with a slight edge over its close competitor, the MaxConf policy.
SelectLLM vs. Baseline Models: The proposed SelectLLM with WeightedMaxConf outperforms individual LLMs baseline by a large margin across both datasets. This indicates that inference with more LLMs boosts performance with a cost of slightly higher latency. Furthermore, the model performs better than the Random, All LLMs and LLM-Blender while being competitive with the Top-$s$ LLMs baseline with respect to accuracy. At the same time, the proposed model has much lower latency (13% lower for GSM8K and 70% lower for MMLU) than Top-$s$ LLMs, indicating the efficiency of the proposed model. This shows that SelectLLM achieves the set goal of better performance than individual LLMs combined with efficiency.
Performance of Oracle vs. other Models: It can be observed that the Oracle score is much higher than that of any individual LLM, baselines, and SelectLLM. This indicates a considerable scope for improvement in the LLM ensembling.
Figure 3: Latency vs. accuracy vs. $s$-values plots for different SelectLLM policies on GSM8K (left) and MMLU (right) datasets. $s$-value: total number of LLMs selected with SelectLLM.
Latency vs. Accuracy vs. $s$-Value: Figure 3 shows the relationship between latencies, accuracies, and $s$-values for different SelectLLM policies across both datasets. WeightedMaxConf consistently outperforms MaxConf and WeightedMaxConf on both accuracy and latency for most $s$-values (except 5 and 6 for GSM8K). This indicates the superiority of WeightedMaxConf, maintaining effectiveness even with a small number of LLMs. A larger number of LLMs and higher latency are required for GSM8K, while a lower $s$-value and latency are needed for MMLU to achieve high accuracy. Selecting 4 LLMs for GSM8K and 2 for MMLU achieves optimal performance, indicating the effectiveness and efficiency of SelectLLM with WeightedMaxConf.
@@
Upper-bound Performance of the SelectLLM: In this section, we investigate potential reasons for the performance gap between the proposed model and the Oracle. With this aim, we estimate the upper-bound performance of SelectLLM with WeightedMaxConf policy. Specifically, we assess all subsets of the top-$s$ LLMs, predicted with SelectLLM for each $q_{i}$. A query is considered solved if at least one subset yields a correct answer. Table 3 presents Oracle, SelectLLM upper bound, and SelectLLM scores across both datasets. The results reveal the following insights: (i) the classifier’s performance is constrained (weighted F1 score for GSM8K is 0.71 and for MMLU is 0.68) due to limited training data (approximately 7K for GSM8K and 14K for MMLU), which likely contributes to the performance gap. Therefore, augmenting the training data or improving the classifier model could enhance scores. (ii) A performance disparity between the best policy and the upper bound suggests the potential for developing better policies (specifically for MMLU). However, since policies rely on the classifier’s confidence scores, enhancing the classifier could also bridge this gap.
Models GSM8K MMLU
Oracle 90.52 90.46
Upper Bound 78.77 76.20
SelectLLM 77.94 65.81
Table 3: Accuracy in different experimental setup.
@@
Ablation Studies: Does the proposed algorithm is effective with different LLMs pool sizes (i.e., value of $k$)? To investigate this question, we conduct ablation studies considering various LLM pool sizes, i.e., $k={1,...,6}$, for the GSM8K dataset. We examine two extreme settings: pool with top-$k$ and bottom-$k$ LLMs based on individual LLM performance. This encompasses many configurations as LLMs with similar or different performances may be present in the LLM pool. We also compare this with closet strong Top-$s$ LLM baseline. The results are shown in the Appendix Figure 7. We observe that even with different $k$ values across both top-$k$ and bottom-$k$ setups, the proposed SelectLLM outperforms (in terms of accuracy) the Top-$s$ LLM baseline. This indicates that the proposed approach is LLM pool-agnostic. Moreover, as the number of $k$ values increases, latency becomes a factor: for larger pool sizes, the latency for SelectLLM is much lower than for Top-$s$ LLM. Similar results are observed with the MMLU dataset.
@@ 7 Conclusions and Future Directions
To the best of our knowledge, this paper presents the first study to efficiently navigate input queries to the most suitable subset of LLMs selected from a large pool. We introduce the novel SelectLLM algorithm, which selects a lightweight subset of LLMs using a multi-label classifier and confidence-based optimal policies. The model is evaluated with two challenging reasoning datasets and compared with several strong baseline models. SelectLLM outperforms individual LLMs and achieves competitive accuracy with a similar subset size of the top-performing LLMs while maintaining significantly lower latency (13% lower latency for GSM8K and 70% lower for MMLU). Although the proposed modeling is promising, we believe better modeling can push the performance closer to the oracle performance. Our findings serve as a strong foundation for such future modeling endeavors. A feasible direction could involve incorporating LLM-related and question-related features to enhance query and LLM awareness in modeling.
@@
One of the key limitations of the proposed SelectLLM algorithm is the limited availability of training data for the multi-label classifier, with only 7K instances for GSM8K and 14K for MMLU. This limitation can potentially lead to biased learning. Despite several measures to address this issue, such as weighing labels to counteract label imbalance, conducting extensive optimal hyperparameter searches, experimenting with different sizes of probabilistic and LLM-based models (with RoBERTa performing the best), and obtaining the best checkpoint with the validation set, the performance remains suboptimal. The algorithm achieves a weighted F1 score of 0.71 for GSM8K and 0.68 for MMLU. Additionally, we could only extract viable answers for 83% to 95% of queries using different LLMs. For the remaining queries, the answers extracted by our algorithm may be invalid or incorrect. These invalid answers could be due to limitations in the extraction algorithm or the LLM itself, where the LLM fails to provide answers in an extractable format.
@@
This paper introduces a novel SelectLLM algorithm to effectively harness the power of LLMs with different capabilities. As the proposed routing models use LLMs, we must acknowledge that, independently of this research, there are certain risks that pertain to all LLMs, as such models may generate outputs that, although plausible, are factually incorrect or nonsensical. Such hallucinations can misguide decision-making and propagate biases, especially in critical scenarios where accuracy is vital. Without proper safeguards, widespread LLM adoption could worsen these concerns. Thus, it is essential to develop mechanisms to mitigate hallucination risks, ensuring responsible and beneficial deployment of these powerful models before adopting the proposed model.
Acknowledgments
We are grateful to the Campus Super Computing Center at MBZUAI for supporting this work.

### v3

@@ 5.1 Baseline Models
Based on recent literature, the following baseline models are included for comparison:
@@
    Random: This represents the mean performance of uniformly randomly selecting an LLM subset from all possible ($12^{k}-$ 1) subsets for each query. We report mean scores across 1,000 independent runs to avoid biases.
@@
    All LLMs: This baseline reports the mean accuracy of maj@(M×$\times|L|$) based on the combined pool of $M$ generations from each LLM and is similar to Li et al. (2024).
@@
    Top-s LLMs: For this baseline, we consider the responses of the top-s scoring LLMs using a majority-vote strategy. The top-performing models are determined by comparing the overall accuracies of the LLMs.
For All LLMs, LLM-Blender and Top-s LLMs, the latency remains constant since they need inference with all LLMs to determine the performance.
@@ 6 Results and Discussion
Top-s LLMs 77.48 19.00 65.75 16.40
SelectLLM MLC + LabelledMaxConf 75.66 14.69 65.68 4.78
MLC + MaxConf 77.48 16.50 65.68 4.78
MLC + WeightedMaxConf 77.94 16.50 65.81 4.78
Table 2: Performance of different models on the GSM8K (Cobbe et al., 2021) and MMLU (Hendrycks et al., 2021) test sets. For GSM8K, we use $s=4$, and for MMLU, $s=2$. The value of $s$ is selected based on experiments on the validation set (see Appendix Section I). Acc: with maj@(M $\times$ $L_{s}$) scores reported in percentage (%); Lat: runtime of $M$ generations for a single query (in seconds); MLC: multi-label classifier; $s$: the number of LLMs considered.
@@
Table 2 presents the performance results for the Oracle, baselines, and the proposed SelectLLM models across both GSM8K and MMLU datasets. We have also reported respective inference latencies to analyze the efficiency of different models. We make the following major observations:
Performance of baselines: We evaluated four baseline models – Random, All LLMs, LLM-Blender, and Top-s LLMs – to assess the impact of ensembling LLMs. The performance of the Random baseline model surpasses that of several individual LLMs reported in Fig. 1, demonstrating the potential of utilizing multiple LLMs. Li et al. (2024) reports that increasing the number of LLMs in the model pool generally increases ensemble performance. Interestingly, however, selecting only the top few LLMs in our experiments performs better than including all LLMs (in all LLMs). Thus, the optimal number of LLMs may depend on the diversity of the model pool (i.e., individual capabilities and overall performance on benchmarks). The Top-s LLMs model proves to be the most effective baseline, outperforming even robust approaches such as LLM-Blender, indicating the benefit of prioritizing the best-performing LLMs. However, the latency of most baseline models is high (as they must utilize the entire model pool at inference), limiting their real-world practicality.
@@
The effect of different policies with the SelectLLM algorithm: It can be observed that the LabelledMaxConf policy yields the lowest performance on the GSM8K dataset. This suboptimal performance may be attributed to the policy’s dependence on both the MLC predictions and confidence scores, wherein the classifier predominantly assigns the label metamath-7b-lm. Given that metamath-7b-lm is a specialized model for mathematics and 88% of the SLData training data is associated with this label, this reliance skews the predictions. However, this effect is minimal for the MMLU dataset, where the label distribution across LLMs is more balanced. This limitation is addressed by the MaxConf and WeightedMaxConf, which relax the constraint on MLC label prediction and only operate on confidence scores. This allows models to incorporate other LLMs and push the performance, particularly for the GSM8K data. Mathematical tuning in WeightedMaxConf allows policies to select LLMs more effectively and improve the scores. Overall, the WeightedMaxConf policy emerged as the best performing, with a slight edge over its closest competitor, the MaxConf policy.
SelectLLM vs. baseline models: The proposed SelectLLM model, utilizing the WeightedMaxConf policy, demonstrates superior performance compared to the Random, All LLMs, and LLM-Blender baselines. Additionally, SelectLLM maintains competitive accuracy relative to the Top-$s$ LLMs baseline while achieving significantly lower latency. Specifically, it reduces latency by 13% for the GSM8K dataset and by 70% for the MMLU dataset compared to all baselines (except Random). This substantial efficiency gain underscores the effectiveness of SelectLLM in balancing performance and operational efficiency.
Figure 3: Quantitative Analysis (GSM8K): Distribution of normalized linguistic feature values (left) and the pairwise absolute difference in normalized feature values between different sets (right).
@@
Upper-bound and performance gap with SelectLLM: We measure the maximum performance achievable by SelectLLM with the WeightedMaxConf policy by establishing an upper-bound performance. Specifically, we compute potential subsets based on the labels predicted by the MLC for each query $q_{i}$. A query is considered solved if at least one subset yields a correct answer. The upper bound scores for GSM8K and MMLU are 78.77 and 76.20, respectively. This reveals the following insights: (1) the classifier’s performance is constrained, with weighted F1 scores of 0.71 for GSM8K and 0.68 for MMLU, due to limited training data (approximately 7K for GSM8K and 14K for MMLU), likely contributing to the performance gap between SelectLLM and Oracle. Therefore, augmenting the training data or improving the classifier model could enhance scores. (2) A performance disparity between the best policy and the upper bound suggests the potential for developing better policies (specifically for MMLU). However, since policies rely on the classifier’s confidence scores, enhancing the classifier could also bridge this gap. We will investigate this performance gap from a linguistic perspective in Section 6.1.
@@
Ablation studies: Is the proposed algorithm effective with different LLMs pool sizes (i.e., value of $k$)? To investigate this question, we conduct ablation studies considering various LLM pool sizes, i.e., $k={1,...,6}$, for the GSM8K dataset. We examine two extreme settings: pools with top-$k$ and bottom-$k$ LLMs based on individual LLM performance. This encompasses many configurations as LLMs with similar or different performances may be present in the LLM pool. We also compare this with closet strong Top-$s$ LLM baseline. The results are shown in the Appendix Figure 8. We observe that even with different $k$ values across both top-$k$ and bottom-$k$ setups, the proposed SelectLLM outperforms (in terms of accuracy) the Top-$s$ LLM baseline. This indicates that the proposed approach is LLM pool-agnostic. Moreover, as the number of $k$ values increases, latency becomes a factor: for larger pool sizes, the latency for SelectLLM is much lower than for Top-$s$ LLM. Similar results are observed with the MMLU dataset.
Set LLM frequency terms time duration units age units small numbers quantifiers ordinals fractional values rates and ratios named entities other units groups
(G1) Solved by All gemma-7b-it -0.938 -0.845 -0.167 -1.0 -0.982 -0.084 -0.138 0.014 1.0 0.073 0.296
gemma-7b-lm -0.936 0.634 0.214 1.0 1.0 0.152 0.743 0.172 1.0 0.699 0.669
llama2-13b-chat -0.988 -0.706 -0.095 -0.999 -0.837 -0.038 -0.264 -0.107 1.0 0.135 0.555
metamath-7b-lm 0.065 0.962 0.33 1.0 1.0 0.279 0.909 0.372 1.0 0.753 0.607
mistral-7b-it -0.993 -0.531 -0.058 -0.842 0.98 -0.023 0.195 0.049 1.0 0.364 0.519
mistral-7b-lm -0.994 -0.452 -0.009 1.0 0.857 -0.009 0.085 -0.002 1.0 0.327 0.472
(G2) Solved by SelectLLM gemma-7b-it -1.0 -0.998 -0.949 -1.0 -1.0 -0.312 -0.996 -0.607 1.0 -0.225 0.911
gemma-7b-lm -1.0 0.775 0.999 1.0 1.0 0.854 -0.268 0.472 1.0 0.913 0.988
llama2-13b-chat -1.0 -0.994 -0.068 -1.0 -1.0 0.021 -0.998 -0.717 1.0 0.081 0.977
metamath-7b-lm 0.45 1.0 1.0 1.0 1.0 0.945 0.995 0.897 1.0 0.964 0.968
mistral-7b-it -1.0 -0.988 0.07 -1.0 1.0 0.34 -0.996 -0.314 1.0 0.504 0.983
mistral-7b-lm -1.0 -0.984 0.448 -1.0 0.989 0.322 -0.994 -0.513 1.0 0.472 0.966
(G3) Solved by Oracle gemma-7b-it -0.674 -0.77 -0.334 -1.0 -0.977 -0.272 -0.176 0.05 -1.0 -0.281 0.037
gemma-7b-lm -0.707 0.185 0.248 0.999 0.902 0.357 0.351 0.332 1.0 0.183 0.129
llama2-13b-chat -0.836 -0.689 -0.183 -1.0 -0.944 -0.116 -0.043 0.042 0.988 -0.241 0.12
metamath-7b-lm 0.102 0.818 0.466 1.0 0.999 0.599 0.62 0.367 1.0 0.475 0.196
mistral-7b-it -0.864 -0.635 -0.194 -1.0 -0.542 -0.027 -0.01 0.195 0.998 -0.197 0.102
mistral-7b-lm -0.88 -0.681 -0.189 -1.0 -0.734 -0.013 0.013 0.115 1.0 -0.166 0.09
(G4) Solved by None gemma-7b-it -0.646 -0.684 -0.817 -1.0 -0.948 -0.279 -0.713 -0.317 -1.0 -0.192 -0.002
gemma-7b-lm -0.629 0.246 0.129 0.997 0.802 -0.07 -0.245 0.171 1.0 0.112 0.198
llama2-13b-chat -0.776 -0.53 -0.742 -1.0 -0.872 -0.259 -0.75 -0.332 -0.981 -0.165 0.113
metamath-7b-lm 0.145 0.763 0.857 1.0 0.992 0.264 0.525 0.513 1.0 0.304 0.304
mistral-7b-it -0.858 -0.508 -0.744 -1.0 -0.463 -0.265 -0.735 -0.132 -0.884 -0.127 0.111
mistral-7b-lm -0.849 -0.47 -0.753 -1.0 -0.606 -0.223 -0.712 -0.237 -0.06 -0.117 0.103
Table 3: Normalized SHAP values for the SelectLLM classifier (GSM8K) over questions from the four subsets of the Test split (see Section 6.1). Each column represents a related group of Synsets. A negative SHAP value (in red) is detrimental to the predicted solvability of corresponding questions, and a positive value (in green) is beneficial.
6.1 Interpretation of Performance Gaps
We investigate the underlying factors contributing to the performance gap between SelectLLM and the Oracle. To gain insights, we perform both linguistic feature-based quantitative analysis and a SHAP values-based qualitative analysis, examining four subsets (i.e., group) of the test data (numbers are specific to the GSM8K dataset): G1: The set of questions correctly solved by all individual LLMs (949 questions), G2: The set of additional questions solved only by SelectLLM and the Oracle (46 questions), G3: The set of additional questions only solved by the Oracle (168 questions) and G4: The set of questions that could not be solved by any LLM (125 questions).
Quantitative Analysis: We extracted 16 linguistic features specific to input questions, as proposed by Srivatsa and Kochmar (2024) for the GSM8K dataset. In Figure 3, we present the feature value distribution (bar plot) and pair-wise value differences (heatmap) across all four subsets. Major insights are: (1) The feature value distributions for G3 and G4 are notably similar, suggesting that questions only solvable by the Oracle exhibit comparable complexity to those not solved by any LLMs. (2) Features such as length, tree depth, NP count, and argument count emerge as influential across the different subsets, which aligns with the findings of Rabinovich et al. (2023) and Srivatsa and Kochmar (2024). A similar observation holds for the MMLU dataset. Future models should incorporate these features to address the performance gap between SelectLLM and the Oracle.
Qualitative Analysis: We extracted word-level SHAP values [^1] Some samples of interactive SHAP plots are hosted at https://anonymous.4open.science/r/SelectLLM-3621. from the SelectLLM classifier for GSM8K. After mapping key content words to WordNet Miller (1994) synsets, we identified synset groups most detrimental to the solvability of each LLM. Table 3 shows the corresponding SHAP values for GSM8K. See Appendix M for more details. We find that: (1) Not all synset groups affect LLMs uniformly; some are detrimental to one model but beneficial to another (e.g., small numbers & quantifiers). Frequency terms (e.g., ‘‘twice", ‘‘thrice") consistently reduce solvability for most LLMs, but MetaMath remains resilient. (2) Gain using SelectLLM: SelectLLM solves questions with more challenging fractional values and rates and ratios than individual LLMs. (3) Gap between SelectLLM and Oracle: quantifiers, age units, other units, and questions about groups like "family" or "team" relatively increase difficulty in G3 compared to G2. (4) Solved by None: G3/G4 rows show that unsolvable questions contain challenging age units, ordinals (e.g., "first", "third"), and named entities.
6.2 Out-of-domain (OOD) Generalization
This section presents a preliminary experiment to validate out-of-domain (OOD) generalization of SelectLLM using the MMLU dataset. As described in Section 4.1, the MMLU dataset comprises 57 category-specific subsets based on grade level (high school and college), subject, domain, etc. We conducted two OOD experiments: one at the grade level and another at the subject level. For the grade-level experiment, the model was trained on high school data and evaluated on college-level data. For the subject-level experiment, we created two non-overlapping splits for training and testing. The results, presented in Table 4, indicate that OOD performance is comparable to in-domain performance, with subject-level generalization yielding even higher accuracy.
Setup Acc ()($\uparrow$)( ↑ ) Lat ()($\downarrow$)( ↓ )
OOD (Grade-level) 58.97 4.80
OOD (Subject-level) 79.75 4.89
In-domain 65.81 4.78
Table 4: OOD experiment results with MMLU dataset for SelectLLM with WeightedMaxConf policy and s=2.
@@ 7 Conclusions and Future Directions
In this work, we introduce an efficient and novel LLM selection algorithm, SelectLLM, to navigate input queries to the most suitable subset of LLMs from a large pool. SelectLLM employs a multi-label classifier and confidence-based optimal policies to select a lightweight subset of LLMs. The model is evaluated on two challenging reasoning datasets and compared against several strong baseline models. SelectLLM demonstrates superior performance compared to these baseline models, achieving competitive accuracy with a similar subset size of top-performing LLMs, while maintaining significantly lower latency. Despite the promising results, we recognize that further advancements in modeling could bring performance closer to the Oracle benchmark. Our findings provide a robust foundation for future research. A potential avenue for improvement could involve integrating LLM-related and query-related features to enhance the model’s query and LLM awareness.
@@
We believe that this study provides a useful starting point for optimal subset selection of LLMs from a large pool. However, we acknowledge that there are certain limitations of this work, and addressing these limitations in the future is an important task.
Extension to general generation tasks
The scope of the current study is limited to two challenging reasoning benchmarks: GSM8K and MMLU. These tasks are framed as generative tasks, where LLMs are required to produce discrete final answers. The study does not focus on general generation tasks such as machine translation, question generation, etc.
It is well known that the one model fits all approach is not universally applicable Huang et al. (2024). For instance, advanced LLMs such as DeepSeek-R1 DeepSeek-AI et al. (2025) demonstrate superior performance in reasoning tasks but exhibit comparatively lower effectiveness in general generation tasks Mercer et al. (2025). Similar observations hold for OpenAI’s and Google’s reasoning models. Building on this observation, we posit that SelectLLM serves as a pivotal step toward the efficient selection of an optimal subset of LLMs from a larger pool, balancing both latency and accuracy, with a focus on reasoning tasks. While the current study does not cover general generation tasks, which are left for future work, the primary challenge for such adaptation lies in implementing robust voting mechanisms. These can be modeled in a simpler manner, as proposed in the All LLMs Li et al. (2023) baseline, to ensure feasibility.
Collection of larger training data for multi-label classifier
Another limitation of the proposed SelectLLM algorithm is the limited availability of training data for the multi-label classifier, with only 7K instances for GSM8K and 14K for MMLU. This limitation can potentially lead to biased learning. Despite several measures to address this issue, such as weighing labels to counteract label imbalance, conducting extensive optimal hyperparameter searches, experimenting with different sizes of probabilistic and LLM-based models (with RoBERTa performing the best), and obtaining the best checkpoint with the validation set, the performance remains suboptimal. The algorithm achieves a weighted F1 score of 0.71 for GSM8K and 0.68 for MMLU.
Reduction in invalid answers
We are able to extract viable answers for 92% to 95% of queries across different LLMs, while the remaining queries resulted in invalid/incorrect outputs from the extraction algorithm. These invalid responses can be attributed to two primary factors: (i) limitations in the LLMs’ response generation, where outputs are not structured in an extractable format, and (ii) limitations in the extraction algorithm, which fails to accurately parse the generated text. Similar limitations have also been reported in previous studies Singh et al. (2023); Chen et al. (2022), particularly for tasks requiring discrete outputs from LLMs. Since the extraction algorithm is the same across all LLMs, the findings are expected to remain the same even if a more effective extraction algorithm is proposed. Future work should prioritize the development of advanced prompting methodologies and optimized LLM architectures that facilitate the generation of discrete, extraction-friendly outputs. Additionally, refining extraction algorithms to improve robustness and adaptability can further reduce the rate of invalid responses.
@@
This paper introduces the SelectLLM algorithm, a novel approach designed to leverage the diverse capabilities of various LLMs. While SelectLLM utilizes LLMs, it is crucial to recognize that, independent of this study, LLMs inherently present risks. These models may generate outputs that, despite being plausible, are factually inaccurate or nonsensical. Such hallucinations can lead to misguided decision-making and the propagation of biases, particularly in high-stakes contexts where accuracy is paramount. In the absence of appropriate safeguards, the broad deployment of LLMs could exacerbate these issues. Thus, it is imperative to develop mechanisms that mitigate the risks of hallucinations to ensure the responsible and effective application of these models.